OpenStudy (anonymous):

Simplify i^72 i -1 -i 1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

same idea as before

OpenStudy (anonymous):

the pattern is \[i^1=i\\ i^2=-1\\ i^3=-i\\ i^4=1\] etc

OpenStudy (anonymous):

so do you divide 72 by 4 again..

OpenStudy (anonymous):

take the integer remainder when you divide \(72\) by \(4\) since \(4\) goes in to \(72\) evenly there is no remainder that makes \[i^{72}=i^0=1\]

OpenStudy (anonymous):

yeah same process every time

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank youu!!

OpenStudy (anonymous):

\[i^{105}=i^1=i\] for example

OpenStudy (anonymous):

yw

jimthompson5910 (jim_thompson5910):

Going the long route \[\Large i^{72} = i^{4*18}\] \[\Large i^{72} = (i^{4})^{18}\] \[\Large i^{72} = (1)^{18}\] \[\Large i^{72} = 1\] ---------------------------------------------------------------- Or the short way, 72/4 = 18 remainder 0 So that means \[\Large i^{72} = i^{0}\] notice how the remainder of 0 is present in the exponent, so, \[\Large i^{72} = i^{0}\] \[\Large i^{72} = 1\] which is a nice shortcut